OpenStudy (anonymous):

a line segment has (x1,y1) as one endpoint and (xm,ym) as its midpoint. Find the other endpoint (x2,y2)of the line segment in terms of x1,y1,xm, and ym.

OpenStudy (koikkara):

SAMPLE A line segment has (x1, y1)as one endpoint and (xm, ym)as its midpoint. Use the formula? (x2, y2) = (2xm − x1, 2ym − y1) to find the coordinates of the endpoint of a line segment if the coordinates of the other endpoint and midpoint are (1, 8), (−3, −1) and (−7, 16), (3, 6),respectively. (a) Endpoint: (1, 8), (−3, −1) (x2, y2) = ( , ) (b) Midpoint: (−7, 16), (3, 6) (x2, y2) = ( , ) ANSWER a) (x2,y2) = (2x-x1, 2y-y1) for midpoints m = 1 = {(2(-3)-1), (2(-1)-8)} = (-7,-10).............................Ans b) (x2,y2) = (2x-x1, 2y-y1) for midpoints m = 1 = {(2(3)+7), (2(6)+16)} = (13,28).............................Ans

OpenStudy (anonymous):

Thank You :)